ROTATIONAL SPECTRA AND STRUCTURE OF PHENYLACETYLENE-H_{2}S COMPLEX

MAUSUMI GOSWAMI and \underline{E. ARUNAN}, Department of Inorganic and Physical Chemistry, Indian Institute of Science, Bangalore 560012, India.

Rotational Spectra of phenylacetylene-H_{2}S, phenylacetylene-HDS, phenylacetylene-D_{2}S, phenylacetylene-H_{2}^{34}S, C_{8}H_{5}D − H_{2}S have been measured. The spectra were recorded using a Balle-Flygare type PNFT Microwave spectrometer[1]. Helium was used as a carrier gas for these measurements. Both ‘a’ and ‘b’ dipole transitions have been observed. For the parent isotopomer, all the transitions are split into two. The rotational spectra of phenylacetylene-HDS indicate that the splitting is due to the interchange of equivalent hydrogens of H_{2}S unit in the complex. The measured rotational constants of the parent species are A = 1206.551(7)MHz, B = 1134.152(6)MHz, C = 732.192(6)MHz for the stronger series and A = 1206.732(7)MHz, B = 1134.056(6)MHz, C = 732.141(8)MHz for the weaker series. The constants derived from the experiments confirm a structure where H_{2}S is placed perpendicular to the plane of the phenylacetylene molecule. This study reveals that the binding of H_{2}O and H_{2}S to the phenylacetylene molecule is very different. Microwave spectroscopic investigations confirmed that H_{2}O prefers to be in plane of the phenylacetylene, donating one of its hydrogen to the acetylenic π cloud while the oxygen of H_{2}O is involved in a secondary interaction forming C-H—O hydrogen bond with the ring hydrogen ortho to the acetylenic group[2].

1. Arunan et al.Current Science,82(2002)533.

2. TA04,63rd International Symposium on Molecular Spectroscopy,Columbus, June 16-20, 2008.